OpenStudy (anonymous):

the sum of two number is 8 find the minumum value of the sum of their squares. a.32 b.16 c.64 d.24

OpenStudy (anonymous):

64-4=60 ans =60 think so

OpenStudy (agent0smith):

Let's call the numbers a and b, and the sum of their squares S: a+b=8 Now get b in terms of a, since we'll need that: b = 8-a \[S = a ^{2}+b ^{2}\] now replace b with (8-a) \[S = a ^{2}+(8-a)^{2}\]

OpenStudy (anonymous):

@agent0smith apply maxima minima

OpenStudy (anonymous):

i think the answer is 32 what do you think guys?

OpenStudy (agent0smith):

To find the minimum value of the sum of their squares, we need to differentiate S w.r.t a: \[S= 2a+2(8-a)(-1)\] And set that equal to zero, find a, and substitute back into the S equation.

OpenStudy (anonymous):

yes the same thing

OpenStudy (anonymous):

oh no it is 56 may be m nt sure

OpenStudy (agent0smith):

Oh and the S in my last reply should be S' not just S. And yep, the answer is 32. You'll find a=4 from the S' = 0 equation, then you put that a value into the equation for S.

OpenStudy (anonymous):

bingo !

OpenStudy (anonymous):

u can even apply AM GM enequality

OpenStudy (anonymous):

You can construct the equation. y = 8- x We can then construct another for the sum of the squares. z = y^2 + x^2 Replace y with x or x with y (whichever you prefer). z = (8-x)^2 + x^2 Then use differentiation to find the minimum point of the curve.

OpenStudy (anonymous):

it 32 for sure

OpenStudy (agent0smith):

^yep, exactly Stephanie.

OpenStudy (anonymous):

thank u guys

OpenStudy (anonymous):

i just checked it.....if you want me to prove??

OpenStudy (anonymous):

no what im saying is u can even apply AM GM enequality